OpenStudy (anonymous):

the speed of train a is 8 mph slower than the speed of train b. train a travels 190 miles in the same time it takes train b to travel 230 miles. find the speed of each train

OpenStudy (anonymous):

come on, there is a question like that everyone hour pretty much

OpenStudy (anonymous):

simultaneous eqns, etc, its boring

OpenStudy (anonymous):

just look at the past hours posts, I gaurantee u there is a similar question

OpenStudy (anonymous):

x and x+8 190/x=230/(X+8) now do it

OpenStudy (anonymous):

40x=190*8 x=38 train a=38 train b=38+8=46